WAEC Mathematics 1992 Question 48

Question 48

Probability

Three balls are drawn one after the other with replacement, from a bag containing 5 red, 9 white and 4 blue identical balls. What is the probability that they are one red, one white and one blue?

  1. A.5/102
  2. B.5/136
  3. C.5/162
  4. D.5/204
  5. E.5/243
Show the answer and explanation

Answer: C (5/162)

Explanation

There are 18 balls. Taking the draws in the order red, white, blue: 5/18 × 9/18 × 4/18 = 180/5832 = 5/162.